# Euler characteristics of the universal Picard stack

### v1

@@ 3.1. The fundamental transformation
$\sum_{n\geq 0}h_{n}=\exp\left(\sum_{n\geq 1}\frac{p_{n}}{n}\right)$
@@
$\mathcal{T}(p_{k})=(p_{k}+1)(1-x^{k})^{-1}-1=\frac{p_{k}+x^{k}}{1-x^{k}},$
@@
$\mathcal{T}(\mathsf{Z}_{g}^{k})=\sum_{m,n\geq 0}\chi^{\mathbb{S}_{n}}_{k}(\mathcal{S}_{g,n}^{m})x^{m}$
@@
$\sum_{n\geq 0}\chi_{k}^{\mathbb{S}_{n}}(\mathcal{S}_{g,n}^{m})=\sum_{n\geq 0}\sum_{i=0}^{m-g}\chi_{k-2i}^{\mathbb{S}_{n}}(\mathrm{Pic}_{g,n})=\sum_{i=0}^{\left\lfloor k/2\right\rfloor}\mathsf{J}_{g}^{k-2i}$
@@
The generating function $\mathsf{J}_{g}^{0}$ is obtained from the expression for $\mathsf{Z}_{g}^{0}$ in [5, Theorem 1.1] by restricting the sum to those Laurent monomials in the $P_{j}$ whose exponents sum to $1$, and then making the substitution $P_{j}\mapsto P_{j}/j$ for all $j$. Specifically,
@@
where the sum is over integers $k,m,s>0$ and $s$-tuples of positive integers $a=(a_{1},\ldots,a_{s})$ and $d=(d_{1},\ldots,d_{s})$, such that
$0<d_{1}<\cdots<d_{s}<m\quad\mbox{and}\quad d_{i}\mid m;$
$\gcd(d_{1},\ldots,d_{s})=1;$
$a_{1}+\cdots+a_{s}=k+1;$
$a_{1}d_{1}+\cdots+a_{s}d_{s}+g-1=km.$
@@
We are left to prove Theorem 3 and Corollary 4.
@@
The generating function $\mathsf{J}_{g}^{top}$ is obtained from Gorsky’s formula [14, Theorem 1] for $\mathsf{Z}_{g}^{top}$ by restricting the sum to Laurent monomials in the $P_{j}$ whose exponents sum to $2$, and then making the substitution $P_{j}\mapsto P_{j}/j$ for all $j$. Explicitly,
@@
Note that Corollary 4 follows immediately from Theorem 3 by setting all of the $P_{i}$ to $1$, which amounts to extracting the degree 0 part of the generating function.
@@
There are small differences between our formula and Gorsky’s [14, Theorem 3.8]. There is one very minor typo in Gorsky’s formula: the definition of $h$ should read
@@
and $k$ should read as $r$. After these substitutions, our formula is extracted directly from his by the transformation outlined in Theorem 3.

### v2

@@ 3.1. The fundamental transformation
$\sum_{n\geq 0}h_{n}=\exp\left(\sum_{n>0}\frac{p_{n}}{n}\right)$
@@
$\mathrm{inv}\Delta(p_{k}\circ\operatorname{Exp}(p_{1}))\circ\operatorname{Log}(1+p_{1})=(p_{k}+1)(1-x^{k})^{-1}-1=\frac{p_{k}+x^{k}}{1-x^{k}}=\mathcal{T}(p_{k}),$
@@
$\mathcal{T}(\mathsf{Z}_{g}^{k})=\sum_{m,n}\chi^{\mathbb{S}_{n}}_{k}(\mathcal{S}_{g,n}^{m})x^{m}$
@@
$\sum_{m,n}\chi_{k}^{\mathbb{S}_{n}}(\mathcal{S}_{g,n}^{m})=\sum_{n}\sum_{i=0}^{m-g}\chi_{k-2i}^{\mathbb{S}_{n}}(\mathrm{Pic}_{g,n})=\sum_{i=0}^{\left\lfloor k/2\right\rfloor}\mathsf{J}_{g}^{k-2i}$
@@
Suppose $g\geq 2$. The generating function $\mathsf{J}_{g}^{0}$ is obtained from the expression for $\mathsf{Z}_{g}^{0}$ in [5, Theorem 1.1] by restricting the sum to those Laurent monomials in the $P_{j}$ whose exponents sum to $1$, and then making the substitution $P_{j}\mapsto P_{j}/j$ for all $j$. Specifically,
@@
where the sum is over integers $k,m,s>0$ and $s$-tuples of positive integers $a=(a_{1},\ldots,a_{s})$ and $d=(d_{1},\ldots,d_{s})$ with $\gcd(d_{1},\ldots,d_{s})=1$ and $d_{i}\mid m$ for each $i$, such that $0<d_{1}<\cdots<d_{s}<m$, $\sum_{i}a_{i}=k+1$, and $\sum_{i}a_{i}d_{i}=km+1-g$.
@@
We are left to prove Theorem 3.
@@
Suppose $g\geq 2$. The generating function $\mathsf{J}_{g}^{top}$ is obtained from Gorsky’s formula [14, Theorem 3.8] for $\mathsf{Z}_{g}^{top}$ by restricting the sum to Laurent monomials in the $P_{j}$ whose exponents sum to $2$, and then making the substitution $P_{j}\mapsto P_{j}/j$ for all $j$. Explicitly,
@@
There are superficial differences between our formula and Gorsky’s [14, Theorem 3.8]. There is one very minor typo in Gorsky’s formula: the definition of $h$ should read
@@
and $k$ should read as $r$. After these substitutions, our formula is extracted directly from his by the transformation defined in Theorem 3.